OpenStudy (anonymous):

Please help :( Given a soda can with a volume of 15 and a diameter of 2, what is the volume of a cone that fits perfectly inside the soda can?

OpenStudy (anonymous):

@hartnn can you help me with this please?

OpenStudy (anonymous):

for this you should have to know what is the volume of can...??? which is like a cylinder..??

OpenStudy (anonymous):

so volume of a cylinder is V=πr^2h

OpenStudy (akashdeepdeb):

So we know the soda can is a cylinder. Volume of cylinder = 15 Diameter = 2 Radius = 1 [Because Diameter = 2 * radius] So, what is the height of the cylinder? Well, we can get an equation! What is the volume of the cylinder? \[\pi * r^2 * h = 15\] r = 1 So, \[\pi * h = 15\] \[h = \frac{15}{\pi}\]  So, the volume of the cone is given by \[\frac{1}{3} \pi * r^2 * h\] Which is \[\frac{1}{3} (The ~~volume~~of~~the~~cylinder!)\] What is the volume of the cylinder? = 15 What is the volume of the cone then? PS The calculation above was not required as we just saw :P Are you getting this? :)

OpenStudy (anonymous):

so the volume of a cone is..?????so you know..????

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

so 1/3 times 15?

OpenStudy (anonymous):

Wow akash has the perfect drawing

OpenStudy (anonymous):

Yes @melanielovee4 :) volume = 1/3 times 15

OpenStudy (anonymous):

basically, divide the volume of the cylinder by 3

OpenStudy (anonymous):

so 5?

OpenStudy (anonymous):

yeeepp..

OpenStudy (anonymous):

ugh i love you guys so much thank you!!

OpenStudy (anonymous):

same to you..!!!